Please solve the following question:
 
The sum 20C0 + 20C1 + 20C2 + …. + 20C10 is:

Grade:11

1 Answers

Meet
6 years ago
As we know that nC0+ nC1+.....+nCn=2^n, and nCr=nC(n-r), so by this formula we can easily write 2[20C0+20C1+....+20C9] +20C10= 2^20 so By this u can easily calculate the value and it comes out to be 2^19+ [(20C10)/2] and its value is 616666